Suppose that two point chargers each with a charge of 1 coulomb are separated by A distance of 6.44 m determine the magnitude of the electrical force between them
1
Expert's answer
2020-04-27T10:16:12-0400
"F=\frac{kq^2}{r^2}"

"F=\frac{8.99\cdot 10^9(1)^2}{6.44^2}=2.17\cdot 10^8\ N"